Wells Technical Institute (WTI), a school owned by Tristana Wells, provides training to individuals who pay tuition directly to the school. WTI also offers training to groups in off-site locations. Its unadjusted trial balance as of December 31, 2018, follows. WTI initially records prepaid expenses and unearned revenues in balance sheet accounts. Descriptions of items a through h that require adjusting entries on December 31, 2018, follow.
  
Additional Information Items

  1. An analysis of WTI's insurance policies shows that $3,864 of coverage has expired.
  2. An inventory count shows that teaching supplies costing $3,349 are available at year-end 2018.
  3. Annual depreciation on the equipment is $15,458.
  4. Annual depreciation on the professional library is $7,729.
  5. On November 1, WTI agreed to do a special six-month course (starting immediately) for a client. The contract calls for a monthly fee of $2,300, and the client paid the first five months' fees in advance. When the cash was received, the Unearned Training Fees account was credited. The fee for the sixth month will be recorded when it is collected in 2019.
  6. On October 15, WTI agreed to teach a four-month class (beginning immediately) for an individual for $5,300 tuition per month payable at the end of the class. The class started on October 15, but no payment has yet been received. (WTI's accruals are applied to the nearest half-month; for example, October recognizes one-half month accrual.)
  7. WTI's two employees are paid weekly. As of the end of the year, two days' salaries have accrued at the rate of $100 per day for each employee.
  8. The balance in the Prepaid Rent account represents rent for December.
WELLS TECHNICAL INSTITUTE
Unadjusted Trial Balance
December 31, 2018
Debit Credit
Cash $ 26,944
Accounts receivable 0
Teaching supplies 10,362
Prepaid insurance 15,545
Prepaid rent 2,073
Professional library 31,088
Accumulated depreciation—Professional library $ 9,328
Equipment 72,533
Accumulated depreciation—Equipment 16,582
Accounts payable 37,202
Salaries payable 0
Unearned training fees 11,500
Common stock 11,000
Retained earnings, December 31, 2017 54,908
Dividends 41,452
Tuition fees earned 105,701
Training fees earned 39,379
Depreciation expense—Professional library 0
Depreciation expense—Equipment 0
Salaries expense 49,743
Insurance expense 0
Rent expense 22,803
Teaching supplies expense 0
Advertising expense 7,254
Utilities expense 5,803
Totals $ 285,600 $ 285,600

4-a. Prepare Wells Technical Institute's income statement for the year 2018.
4-b. Prepare Wells Technical Institute's statement of retained earnings for the year 2018.
4-c. Prepare Wells Technical Institute's balance sheet as of December 31, 2018.
Answer #1

4-a. Prepare Wells Technical Institute's income statement for the year 2018.

Wells Technical Institute (WTI)
Income Statement
For the Year Ended December 31, 2018
Revenue:
Tuition fees earned $118951
Training fees earned 43979
Total revenues 162930
Expenses:
Depreciation expense—Professional library 7729
Depreciation expense—Equipment 15458
Salaries expense 50143
Insurance expense 3864
Rent expense 24876
Teaching supplies expense 7013
Advertising expense 7254
Utilities expense 5803
Total expenses 122140
Net income $40790

Tuition fees earned= $105701+($5300*2.5)= $118951

Training fees earned= $39379+($2300*2)= $43979

Salary expense= $49743+($100*2*2)= $50143

Rent expense= $22803+2073= $24876

Teaching supplies expense= $10362-3349 = $7013

b. Prepare Wells Technical Institute's statement of retained earnings for the year 2018.

Wells Technical Institute (WTI)
Statement of Retained Earnings
For the Year Ended December 31, 2018
Retained earnings, December 31, 2017 $54908
Add: Net income 40790
95698
Less: Dividends 41452
Retained earnings, December 31, 2018 $54246

c. Prepare Wells Technical Institute's balance sheet as of December 31, 2018.

Wells Technical Institute (WTI)
Balance Sheet
For the Year Ended December 31, 2018
ASSETS
Cash $26944
Accounts receivable 13250
Teaching supplies 3349
Prepaid insurance 11681
Professional library 31088
Accumulated depreciation—Professional library 17057 14031
Equipment 72533
Accumulated depreciation—Equipment 32040 40493
Total assets $109748
LIABILITIES
Accounts payable 37202
Salaries payable 400
Unearned training fees 6900
Total liabilities 44502
STOCKHOLDERS' EQUITY
Common stock 11000
Retained earnings 54246 65246
Total liabilities and stockholders' equity $109748

Accounts receivable= $0+(5300*2.5)= $13250

Prepaid insurance= $15545-3864= $11681

Accumulated depreciation—Professional library= $9328+7729= $17057

Accumulated depreciation—Equipment= $16582+15458= $32040

Salaries payable= $100*2*2= $400

Unearned training fees= $11500-(2300*2)= $6900